I'm having trouble will a few similar questions. I think once I see this one done, I will be able to get the rest. I need the point slope form of the line's equation. Then using the point-slope form, I need t o write the slope-intercept in function notation.
Given conditions: Slope = 5, passing through (2,3)
Thank you so much!

Answer by John10(245)

Okay...point-slope formula: y - yo = m(x - xo) then simplify to get slope-intercept formula as y = mx + b where (xo,yo) is the given point and m is the slope.
John10:)